OpenStudy (anonymous):

find the equations for the asymptotes of the hyperbola (x^2)/36 - (y^2)/5 = 1 please explain in detail?

OpenStudy (amistre64):

the bottom give you the measure of a^2 and b^2 in a standard hyper notation

OpenStudy (amistre64):

the slope of the asymptote is y/x, or in this case sqrt(5)/sqrt(36) and it goes thru the point (0,0) since its centerd at zero

OpenStudy (amistre64):

x^2/a^2 - y^2/b^2 = 1 (x/a)^2 - (y/b)^2 = 1 (under y)/(under x) = slope right?

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=%28x^2%29%2F36+-+%28y^2%29%2F5+%3D+1 click on "properties" for a dbl check

OpenStudy (anonymous):

lol im still trying 2 unerstand but thnx 4 th help

OpenStudy (amistre64):

practice helps :)